evaluate ∫ 1 / 1 + cos 2x dx

evaluate
∫ 1 / 1 + cos 2x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{1}{1+cos2x}dx
I = \int \frac{1}{2cos^2x}dx
I = \int \frac{sec^2x}{2}dx
I = \frac{tanx}{2}+c